# Can Large Language Models Generate High-quality Patent Claims?

### v1

@@ 1 Introduction
Patent claims are the critical part of a patent application that defines the boundaries of patent protection. Drafting patent claims necessitates the expertise of qualified patent agents or attorneys. It requires a profound understanding of the invention’s technical details and a familiarity with pertinent patent laws and language conventions (Faber, 1990). Meticulous and precise patent claims are essential for securing robust patent protection. However, drafting and revising patent applications are both time-intensive and financially demanding. Each round of revising can generate costs up to thousands of dollars (LLP, 2023), presenting considerable challenges particularly for small enterprises. Therefore, automation through appropriate assistive language models could significantly enhance quality and efficiency.
As shown in Figure 1, previous studies have mainly focused on generating claims based on patent abstracts (Lee, 2020). However, the performance of these methods is not satisfying and the task appears ill-posed as abstracts do not fully disclose all details of an invention. In contrast, patent descriptions provide detailed explanations of the invention and its operation. It must, by law, fully disclose the details that the claims summarize and condense into the gist (European Patent Office, 2000). Therefore, this paper investigates description-based claim generation and compares it with previous solely abstract-based studies.
@@
Given the significant difficulty of the task, the purpose of this research is to evaluate the performance of current LLMs for description-based claim generation. We aim to investigate the following research questions (RQs): RQ1: Does description-based claim generation outperform abstract-based generation? RQ2: Does current patent-specific LLM outperform general LLMs on claim generation? RQ3: On which aspects can current LLMs do well already and on which cannot?
The main contributions are as follows.
1. We construct the first dataset for description-based claim generation, namely HUPD-DCG. We demonstrate theoretically and empirically that description-based claim generation outperforms previous abstract-based generation, particularly increasing the invention’s feature completeness.
2 Background and Related Work
2.1 Patent description, abstract, and claims
@@
We present an example patent description, abstract, and claims in Appendix A. The patent description provides a comprehensive illustration of the invention, including structural and functional properties of the invention based on detailed embodiments and examples as well as the context in which it is used. The abstract is a short summary that provides a quick overview of the technical invention in the patent document. It is usually limited to a few sentences and must be concise, covering the main technical aspects without going into details. Patent claims are the legally significant part of a patent application, defining the technical scope of the invention to withstand legal scrutiny and secure legal protection. Claims must be crafted with high-level precision and clarity, comprising technological aspects that warrant protection and include the features that constitute the invention.
2.2 Patent Text Generation
Jiang and Goetz (2024) summarized four types of text generation tasks in the patent field, including summarization, translation, simplification, and patent writing. We list open-sourced datasets aiming at patent text generation in Table 2. Patent summarization (Sharma et al., 2019) and translation (Pouliquen, 2015) have been extensively studied. Notably, patent translation systems are provided by various patent offices, including the European Patent Office’s Patent Translate.222 https://www.epo.org/en/searching-for-patents/helpful-resources/patent-translate
@@ 3.1 Dataset
In this research, we construct the first dataset for description-based claim generation. The Harvard USPTO Patent Dataset (HUPD) Suzgun et al. (2024) is a recently collected large-scale multi-purpose patent dataset, including more than 4.5 million patent documents with 34 data fields (patent description, abstract, and claims are included). [^3] The USPTO stands for the United States Patent and Trademark Office, granting U.S. patents written in English. https://www.uspto.gov/ We propose our dataset, HUPD-DCG (Description-based Claim Generation), by filtering a portion of target documents based on this large dataset.
Firstly, we selected all the patent documents filed in 2017. Then, we eliminated any pending or rejected documents and only kept granted documents to formulate a high-quality dataset for claim generation. Considering the context length of Llama-3, we opted for documents with a description length smaller than 8,000 tokens in our experiments. In practical settings, models are developed by training on existing patent documents and subsequently employed for new applications. To simulate realistic scenarios, we ordered the documents by date and used the last 1,311 documents for testing, which is about 14% of the whole dataset.
@@ 3.2 Models
GPT-4 We use the latest GPT-4 model (OpenAI, 2023), representing the state-of-the-art LLM with substantial general knowledge and outstanding reasoning capabilities.444gpt-4-turbo-2024-04-09: https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo Notably, GPT-4 shows remarkable performance on summarization tasks (Pu et al., 2023). Hence, we do not evaluate other LLMs specifically trained for summarization.
Llama-3 Among open-source LLMs, we choose the most recent Llama-3, pre-trained on over 15 trillion tokens of data from publicly available sources. [^5] https://llama.meta.com/llama3/ Llama-3 outperforms most open-source models on common industry benchmarks. We opt for the Llama-3-8B-Instruct version because it is less computationally intensive and has been fine-tuned to execute user instructions more accurately and effectively.
Llama-3-FT Llama-3-FT is a fine-tuned version of the original Llama-3-8B-Instruct model based on our train set. We use LoRA (Hu et al., 2021) for training, a parameter-efficient approach to reduce computational needs while maintaining comparable performance. The inputs are instruction prompts and patent descriptions, while the output is corresponding patent claims. Appendix B lists experimental details.
Llama-3-MFT Multi-task fine-tuning (MFT) has shown effectiveness in many applications, such as code generation (Liu et al., 2023). MFT trains the model on multiple tasks simultaneously to enhance its ability to better generalize across a broader range of tasks by learning shared and task-specific features. Our Llama-3-MFT model is trained on two tasks, description-based claims generation and summarization (abstract generation). Both tasks require processing complex documents and generating shorter texts. Thus, we aim to investigate whether MFT could bring extra benefits to claim generation compared to single fine-tuning.
@@ 3.3 Evaluation Metrics
Human evaluation Human evaluation for patent claims needs substantial patent expertise. However, involving patent experts to evaluate large numbers of claim sets is prohibitively costly and labor-intensive. [^6] Hourly rates per attorney are about \$350 to \$750. Therefore, we specifically focus on ten documents and conduct a comparative study of multiple versions of patent claims generated by Llama-3, Llama-3-FT, Llama-3-MFT, and GPT-4. A licensed patent attorney and an experienced patent engineer, both with extensive experience in drafting patent applications, conduct the evaluation and agree on the results. The comparative analysis aims to highlight the strengths and weaknesses of each LLM in generating patent claims. We also provide insights into the potential of LLMs to automate or partially automate the patent drafting process. Five criteria are employed to evaluate the quality of the patent claims and details of the rating strategies are listed in Appendix C.
1. Completeness of Essential Features: The extent to which the generated claims encapsulated all critical aspects of the invention. 2. Conceptual Clarity: The clarity and unambiguity of the language used in the claims. 3. Consistency in Terminology: The uniformity in the use of terms throughout the claims. 4. Technical Correctness of Feature Linkages: The accuracy with which the features were interconnected and related. 5. Overall Quality: An aggregate measure combining all the above criteria.
@@ 4.1 RQ1: Does description-based claim generation outperform previous abstract-based methods?
Model BLEU R-1 R-L BERTScore
GPT-2-FT (abstract) - 52.38 - -
Llama-3 (abstract) 21.98 46.30 31.38 86.17
Llama-3 (description) 34.32 56.72 38.44 88.40
Table 4: Comparison of abstract-based and description-based claim generation. The description-based method shows higher evaluation scores. Note that the dataset used in the previous study is not publicly available, so the results are not directly comparable.
4.2 RQ2: Does current patent-specific LLM outperform general LLMs?
Table 5 presents a comparison of the performance metrics for five different models. Although PatentGPT-J is specialized for patent texts, it shows significantly lower scores across all metrics. Additionally, our manual examination of the claims generated by PatentGPT-J revealed that they are often nonsensical and lacking in detail. The claims tend to be short and repetitive, further underscoring the model’s limitations.
Takeaways PatentGPT-J demonstrates poor performance on claim generation, indicating that current patent-specific LLMs are far from satisfactory. However, domain-specific LLMs have significant potential to perform in-domain tasks. Therefore, we suggest research groups with enough resources can pre-train such models for this important field.
Automated Evaluation Human Evaluation
Model BLEU R-1 R-L BERTScore Feature Completeness Conceptual Clarity Terminology Consistency Correctness of Feature Linkage Overall Quality
PatentGPT-J 12.86 30.68 23.07 80.24 - - - - -
Llama-3 35.42 57.17 38.88 88.49 4.0 4.8 6.6 4.1 4.6
Llama-3-FT 37.52 59.96 47.35 89.45 5.3 5.8 6.9 5.3 5.6
Llama-3-MFT 37.27 58.86 46.25 89.41 5.1 4.9 6.5 5.3 5.4
GPT-4 15.73 52.59 37.73 87.34 5.4 6.3 7.4 5.9 6.0
Table 5: Automated and human evaluation results. The best scores for each metric are marked in bold. PatentGPT-J shows poor performance and nonsensical outputs. Fine-tuning improves on both automated and human evaluation metrics. Multi-task fine-tuning performs slightly worse than single fine-tuning. GPT-4 demonstrates lower automated evaluation scores but has the best overall quality based on human evaluation.
@@ 4.2 RQ2: Does current patent-specific LLM outperform general LLMs?
4.3 RQ3: On which aspects can current LLMs do well already and on which cannot?
Table 5 shows the human evaluation results of Llama-3, Llama-3-FT, Llama-3-MFT, and GPT-4.
Feature Completeness In most cases, each first claim generated by Llama-3, Llama-3-FT, Llama-3-MFT, and GPT-4 is found to be error-free and accurately reflects the reference claims and the description. This finding aligns with the automated evaluation results listed in Table 6. The scores of the first claim on all metrics are largely higher than the results on the remaining claims. This accuracy is likely because the description often includes the independent claim verbatim before introducing additional optional features and embodiments. Consequently, the model’s ability to extract and replicate this claim correctly is understandable, given the straightforward nature of the task.
@@ 4.3 RQ3: On which aspects can current LLMs do well already and on which cannot?
We find two key improvements of completeness on problems (1) and (2) from Llama-3-FT. Some dependent claims include additional features that provide some level of limitation or specificity to the respective independent claim it referenced. Moreover, Llama-3-FT rarely misclassifies features, representing a significant improvement in adhering to standard practice. Therefore, fine-tuning improves the completeness scores from 4.0 to 5.3, but significant gaps still remain. The feature coverage is not good enough and the fine-tuned models also generate repetitive claims. Human evaluation aligns with automated evaluation, where the fine-tuning of Llama-3 improves on all automated metrics compared to its zero-shot performance, especially about an 8.5% gain in R-L. This suggests fine-tuning leads to more lexical overlaps and semantic similarity with referenced claims.
Additionally, the features claimed in referenced dependent claims are more comprehensively covered by GPT-4 compared to other models with a completeness score of 5.4, although not all features are well represented. Furthermore, GPT-4 avoids outputting repetitions of claim features, which is an improvement compared to Llama-3-based models.
@@
Nonetheless, GPT-4 results in the best human evaluation. The primary claim is mostly accurate, and the dependent claims exhibit better technical relevance, clarity, and logical grouping. The model demonstrates a better understanding of technical features, their interactions, and their significance within the overall context of the invention, indicating its capability to evaluate the entire text and generate a coherent claim set.
@@
Takeaways All models can generate a high-quality first independent claim, but perform poorly on dependent claims. Fine-tuning brings two major improvements in feature completeness, avoiding replicating claims without adding additional features and avoiding misclassifying features. Although MFT keeps the improvement brought by fine-tuning, new problems are introduced, significantly decreasing conceptual clarity. Additionally, claims generated by all models demonstrate relatively high terminology consistency, but further improvements are needed. From the perspective of the correctness of feature linkage, GPT-4 selects and groups features within dependent claims logically, whereas Llama-3 randomly chooses features and describes them separately. Although GPT-4 generates short texts and has less sequence coverage with referenced claims, it achieves a higher rating in all human evaluation metrics. Despite promising, there are challenges unsolved, particularly in achieving the completeness of essential features and maintaining technical accuracy.

### v2

@@ 1 Introduction
Patent claims are the critical part of a patent application that defines the boundaries of patent protection. We introduce the background of patent descriptions, abstracts, and claims in Appendix A. Drafting patent claims necessitates the expertise of qualified patent agents or attorneys. It requires a profound understanding of the invention’s technical details and a familiarity with pertinent patent laws and language conventions (Faber, 1990). Meticulous and precise patent claims are essential for securing robust patent protection. However, drafting and revising patent applications are both time-intensive and financially demanding. Each round of revising can generate costs up to thousands of dollars (LLP, 2023), presenting considerable challenges particularly for small enterprises. Therefore, automation through appropriate assistive language models could significantly enhance quality and efficiency.
As shown in Figure 1, previous studies have mainly focused on generating claims based on patent abstracts for simplicity (Lee, 2020). However, the performance of these methods is not satisfying and the task appears ill-posed as abstracts do not fully disclose all details of an invention. In contrast, patent descriptions provide detailed explanations of the invention and its operation. It must, by law, fully disclose the details that the claims summarize and condense into the gist (European Patent Office, 2000). Therefore, this paper investigates description-based claim generation and compares it with previous solely abstract-based studies.
@@
Since drafting patent claims by qualified patent agents or attorneys is both time-consuming and financially burdensome, we aim to explore whether this process can be realized by LLMs. We design three research questions (RQs) to explore whether LLMs can generate high-quality patent claims.
RQ1: Does description-based claim generation outperform abstract-based generation? In principle, patent claims can be generated from any part of a patent document. Previous studies have focused on using patent abstracts for claim generation, despite their brevity and limited legal significance. Since the description is the legal centerpiece, we proposed claim generation based on the description rather than other parts and compared it with previous abstract-based methods. [^2] It is worth noting that combining abstract and description is neither legally nor technically judicious. Legally: the abstract does not contribute extra information to the description and if it did, it must not be used in the claims (37 US CFR 1.72 and 1.438/PCT Rule 8); the claims have to be solely rooted in the description. Technically: the use of the abstract is rather a wish driven by model limitations; as soon as a model can manage the description, the abstract entirely loses its value as it cannot contain any useful content beyond the description legally anyway; if a model has additional bandwidth for input, there may rather be the question if it could be used for adding external information, e.g., close prior art or similar. Future research should rather study that. Therefore, inputting both abstract and description may be unnecessary and even complicate the task.
RQ2: Does current patent-specific LLM outperform general LLMs on claim generation? It is shown that small-scale domain-specific LLMs (e.g., 7B) can outperform strong general LLMs (e.g., GPT-4) on specific tasks, such as in math reasoning (Shao et al., 2024). Meanwhile, the legal-specific LLM SaulLM-7B is reported to outperform Mistral-7B and Llama-2-13B on LegalBench-Instruct and Legal-MMLU benchmarks (Colombo et al., 2024). Therefore, we included legal and patent-specific LLMs to investigate whether domain-specific LLMs can perform better on this task.
RQ3: How do LLMs perform on patent claims’ completeness, clarity, consistency, and logical linkage? To make comprehensive evaluations, we have tested with different model architectures, different sizes, and models in specific domains. We conduct human evaluations that adhere strictly to established examination criteria to assess LLMs’ capabilities and limitations more accurately.
The main contributions are as follows:
1. We construct the first dataset for description-based claim generation. We demonstrate theoretically and empirically that description-based claim generation outperforms previous abstract-based generation, particularly increasing the invention’s feature completeness.
2 Related Work
Jiang and Goetz (2024) summarized four types of text generation tasks in the patent field, including summarization, translation, simplification, and patent writing. We list open-sourced datasets aiming at patent text generation in Table 2. Patent summarization (Sharma et al., 2019) and translation (Pouliquen, 2015) have been extensively studied. Notably, patent translation systems are provided by various patent offices, including the European Patent Office’s Patent Translate.333 https://www.epo.org/en/searching-for-patents/helpful-resources/patent-translate.
@@ 3.1 Dataset
In this research, we construct the first dataset for description-based claim generation. The Harvard USPTO Patent Dataset (HUPD) Suzgun et al. (2024) is a recently collected large-scale multi-purpose patent dataset, including more than 4.5 million patent documents with 34 data fields (patent description, abstract, and claims are included). [^4] The USPTO stands for the United States Patent and Trademark Office, granting U.S. patents written in English. We propose our dataset, HUPD-DCG (Description-based Claim Generation), by filtering a portion of target documents based on this large dataset.
Firstly, we selected all the patent documents filed in 2017. We eliminated any pending or rejected documents and only kept granted documents to formulate a high-quality dataset for claim generation. Considering the context length of some LLMs, such as Llama-3, we opted for documents with a description length smaller than 8,000 tokens in our experiments. In practical settings, models are developed by training on existing patent documents and subsequently employed for new applications. To simulate realistic scenarios, we ordered the documents by date and used the last 1,311 documents for testing, which is about 14% of the whole dataset.
@@ 3.2 Models
We select the recent Llama-3-8B [^5] https://llama.meta.com/llama3/ and Mistral-7B (Jiang et al., 2023) as the base models because they have shown outstanding capabilities and represent different model structures. To investigate the size effect, we opt for Llama-3-70B, Mixtral-8$\times$7B (Jiang et al., 2024), Mixtral-8$\times$22B, and state-of-the-art GPT-4 (OpenAI, 2023) model. In addition, we conduct experiments on domain-specific LLMs, including the largest open-source patent LLM PatentGPT-J-6B (Lee, 2023), and the recent legal-specific LLM SaulLM-7B (Colombo et al., 2024). Furthermore, we fine-tune the Llama-3-8B model based on our train set. Llama-3-FT is the version of normal single-task fine-tuning, while Llama-3-MFT is trained on two tasks simultaneously using multi-task fine-tuning (MFT), including description-based claim generation and summarization (abstract generation). We introduce details of each model in Appendix B and experimental details in Appendix C.
@@ 3.3 Evaluation Metrics
Human evaluation Although automated metrics provide standardized and reproducible ways to compare model performance based on lexical overlaps or semantic relationships, we suspect them to struggle. We supplement our approach with human evaluations conducted by patent professionals. These evaluations adhere strictly to established examination criteria, enhancing the accuracy of our results. Readers should primarily focus on human evaluation results, whereas automated metrics may provide valuable insights at times.
Human evaluation for patent claims needs substantial patent expertise. However, involving patent experts to evaluate large numbers of claim sets is prohibitively costly and labor-intensive. [^6] Hourly rates per attorney are about \$350 to \$750. Therefore, we specifically conduct the human evaluation on eighty patent claims of multiple versions generated by LLMs. [^7] Since the outputs generated by domain-specific LLMs were mostly not sensible, we did not conduct human evaluations on those results. A licensed patent attorney and an experienced patent engineer, both with extensive experience in drafting patent applications, conduct the evaluation and agree on the results.
Five criteria are employed to evaluate the quality of the patent claims and details of the rating strategies are listed in Appendix E. The criteria include: 1. Completeness of Essential Features: The extent to which the generated claims encapsulated all critical aspects of the invention. 2. Conceptual Clarity: The clarity and unambiguity of the language used in the claims. 3. Consistency in Terminology: The uniformity in the use of terms throughout the claims. 4. Technical Correctness of Feature Linkages: The accuracy with which the features were interconnected and related. 5. Overall Quality: An aggregate measure combining all the above criteria.
@@ 4.1 RQ1: Does description-based claim generation outperform previous abstract-based methods?
Model BLEU R-1 R-L BERTScore
GPT-2-FT (abstract) Lee (2020) - 52.38 - -
Llama-3 (abstract) 21.98 46.30 31.38 86.17
Llama-3 (description) 34.32 56.72 38.44 88.40
Table 4: Comparison of abstract-based and description-based claim generation. Note that the dataset used in the previous study is not publicly available, so we list previous results as a reference, which are not directly comparable to our experiments and results. This table showed that the description-based claim generation outperformed previous abstract-based methods.
@@
Human Evaluation Automated Evaluation
Model Feature Completeness Conceptual Clarity Terminology Consistency Correctness of Feature Linkage Overall Quality BLEU R-1 R-L BERTScore
Domain-specific LLMs
PatentGPT-J-6B - - - - - 12.86 30.68 23.07 80.24
SaulLM-7B - - - - - 12.68 36.63 25.10 83.13
Base LLMs
Mistral-7B 5.0 4.3 4.6 5.0 4.8 29.70 49.17 36.20 85.33
Llama-3-8B 4.0 4.8 6.6 4.1 4.6 35.42 57.17 38.88 88.49
Fine-tuned LLMs
Llama-3-FT 5.3 5.8 6.9 5.3 5.6 37.52 59.96 47.35 89.45
Llama-3-MFT 5.1 4.9 6.5 5.3 5.4 37.27 58.86 46.25 89.41
Large-sized LLMs
Llama-3-70B 5.6 5.0 5.2 5.0 5.3 36.40 59.89 41.79 87.44
Mixtral-8$\times$7B 5.7 5.0 5.0 5.0 5.2 37.03 60.18 43.67 88.51
Mixtral-8$\times$22B 5.6 4.8 5.4 5.2 5.3 33.96 60.57 46.30 88.97
GPT-4 5.4 6.3 7.4 5.9 6.0 15.73 52.59 37.73 87.34
Table 5: Automated and human evaluation results. The best scores for each metric are marked in bold. PatentGPT-J and SaulLM show poor performance and sometimes nonsensical outputs. Fine-tuning improves on both automated and human evaluation metrics. Multi-task fine-tuning performs slightly worse than single fine-tuning. GPT-4 demonstrates lower automated evaluation scores but has the best overall quality based on human evaluation.
@@
4.2 RQ2: Does current patent-specific LLM outperform general LLMs?
Table 5 presents a comparison of the performance metrics for ten different models. Although PatentGPT-J is specialized for patent texts, it shows significantly lower scores across all metrics. Additionally, our manual examination of the claims generated by PatentGPT-J revealed that they are often nonsensical and lacking in detail. The claims tend to be short and repetitive, further underscoring the model’s limitations.
Additionally, SaulLM performs slightly better than PatentGPT-J but still substantially worse than others. Based on our observation, the outputs of SaulLM can be categorized into the following cases: (1) The generated texts have a similar format to patent claims. This indicates that the legal-specific LLM can somehow capture the structure and format of patent claims. Although such outcomes are rare, they demonstrate the potential of legal-specific LLMs on patent claim generation. A possible reason is that SaulLM has been trained on structured patent texts, allowing it to replicate this format in its outputs. (2) The output is a short narrative paragraph of what the patent has claimed. This result shows the model’s tendency towards narrative text instead of structured patent claims. This tendency can be attributed to the model’s training predominantly on narrative and explanatory legal content, rather than strictly formatted patent claims. (3) The model generates a summary of the patent description. It suggests that the model might have misunderstood the input prompts. The instruction-tuning process of SaulLM may not be sufficient or the model tends to recognize patent claims as short summaries. (4) The output makes no sense or repeats the same sentence multiple times. This model’s inability may be due to insufficient training and data. Whereas Llama-3 is pre-trained on 15 trillion tokens, SaulLM is based on about 30 billion tokens.
Takeaways (1) PatentGPT-J demonstrates poor performance on claim generation, indicating that current patent-specific LLMs are far from satisfactory. We suggest research groups with enough resources can pre-train such models for this important field. (2) SaulLM designed for the legal domain also under-performs compared to other general LLMs. Although the patent is one type of legal document, it features specific language and format requirements that differentiate it from others. Hence, LLMs for the general legal domain may not be well-suited for patent claim generation.
4.3 RQ3: How do LLMs perform on patent claims’ completeness, clarity, consistency, and logical linkage?
Feature Completeness In most cases, each first claim generated by LLMs is found to be error-free and accurately reflects the reference claims and the description. This finding aligns with the automated evaluation results listed in Table 6. The scores of the first claim on all metrics are largely higher than the results on the remaining claims. This accuracy is likely because the description often includes the independent claim verbatim before introducing additional optional features and embodiments. Consequently, the model’s ability to extract and replicate this claim correctly is understandable, given the straightforward nature of the task.
@@ 4.3 RQ3: How do LLMs perform on patent claims’ completeness, clarity, consistency, and logical linkage?
We find two key improvements of completeness on problems (1) and (2) from Llama-3-FT. Some dependent claims include additional features that provide some level of limitation or specificity to the respective independent claim it referenced. Moreover, Llama-3-FT rarely misclassifies features, representing a significant improvement in adhering to standard practice. Therefore, fine-tuning improves the completeness scores from 4.0 to 5.3, but significant gaps still remain. The feature coverage is not good enough and the fine-tuned models also generate repetitive claims. Human evaluation aligns with automated evaluation, where the fine-tuning of Llama-3 improves on all automated metrics compared to its zero-shot performance, especially about an 8.5% gain in R-L. This suggests fine-tuning leads to more lexical overlaps and semantic similarity with referenced claims. Meanwhile, larger models can also bring this type of improvement. Llama-3-70B increases the R-L from 38.88% to 41.79% compared to the 8B version, and Mistral-8$\times$22B improves from 43.67% to 46.30% versus its 8$\times$7B version.
Additionally, the features claimed in referenced dependent claims are more comprehensively covered by larger models, such as Mixtral-8$\times$7B with a completeness score of 5.7, although not all features are well represented. Furthermore, GPT-4 avoids outputting repetitions of claim features, which is an improvement compared to Llama-3-based models.
@@
Nonetheless, GPT-4 results in the best human evaluation. The primary claim is mostly accurate, and the dependent claims exhibit better technical relevance, clarity, and logical grouping. The model demonstrates a better understanding of technical features, their interactions, and their significance within the overall context of the invention, indicating its capability to evaluate the entire text and generate a coherent claim set. Due to the large inconsistency between human and existing automatic evaluation methods, a promising research direction is to develop new automated metrics that have closer alignment to human evaluation, such as using LLM-based evaluators (Liu et al., 2023b).
@@
Takeaways (1) All models can generate a high-quality first independent claim, but perform poorly on dependent claims. (2) Fine-tuning brings two major improvements in feature completeness, avoiding replicating claims without adding additional features and avoiding misclassifying features. (3) Although MFT keeps the improvement brought by fine-tuning, new problems are introduced, significantly decreasing conceptual clarity. (4) Claims generated by all models demonstrate relatively high terminology consistency, but further improvements are needed. (5) Larger models can select and group features within dependent claims logically, instead of randomly choosing features and describing them separately. (6) Although GPT-4 generates short texts and has less sequence coverage with referenced claims, it achieves a higher rating in all human evaluation metrics. (7) Despite promising, there are challenges unsolved, particularly in achieving the completeness of essential features and maintaining technical accuracy.